# A Spatial Gap in the Sky Distribution of Fast Radio Burst Detections Coinciding with Galactic Plasma Overdensities

### v1

@@ 4.4 Empirical Limits on Scattering Timescale
As summarized in §4.1, the measurements to date of sources observed through the Cygnus X region – through angular broadening of background quasars and Cygnus X-3, scattering of distant pulsars and one FRB detected by FAST – show significant scattering ($\tau_{\rm sc,\,600\,MHz}\sim 10-1000$ ms). This already suggests that scattering can be deleterious for FRB detections at CHIME, and is likely a significant contributor to the lack of CHIME/FRB detections in the gap. However, the magnitude of scattering, which depends on plasma fluctuations on tiny scales at $\ll\rm{AU}$, can drastically differ between sightlines. In this Section, we quantify the effect of scattering on CHIME/FRB detections, and in turn set a lower limit on the average scattering in the detection gap which would lead to zero detections.
@@
We first searched the positional and fitburst data in Catalog 2 for a radial trend in measured $\tau_{\rm sc}$ from the boundary of the Cygnus X region, observing no clear trend. In Catalog 1 injections, CHIME/FRB found a strong selection bias against temporally broad bursts ($W_{b}>10$ ms, calculated at 600 MHz; M. Merryfield et al. 2023). The lack of a trend may be due to the narrow observable window in $\tau_{\rm sc}$: the lower limit set by the reliability cut-off (see §2.4; 0.13 ms at 1 GHz); and the upper limit ($\geq 10$ ms at 1 GHz) that corresponds to scattered pulses likely too broad to detect. Along with a limited number of detected FRBs, a statistically significant radial trend is difficult to establish without finer time resolution to probe a wider range in $\tau_{\rm sc}$. As a result, it is difficult to trace if there is a gradual increase in scattering timescales for sightlines closer to the center of the detection gap.
@@
From the simulations described in §3.2, we found that on an average 6 bursts are expected to be seen in the detection gap. These simulations account for the elevated sky temperature but not for scattering. Here, we extend the simulation to assess the degree of scattering required to see zero bursts.
@@
Figure 3: Distribution of scattering timescales at 1 GHz that result in zero detectable FRBs in the gap across 106 simulations, incorporating elevated sky temperature as described in §3.2. Vertical dashed lines indicate scattering predictions from the NE2001 model (black) and YMW16 model (red) for the line of sight along the center of the gap.
@@
In the simulations outlined in §3.2, we introduce pulse broadening due to scattering using Equation 5. For each simulated realization, we compute the minimum $\tau_{\rm sc}$ using Equation 4, such that the increased $W_{b}$ and $T_{\rm sky}$ reduce the signal-to-noise of all FRBs within the detection gap to $(S/N)_{b}<(S/N)_{\rm thresh}=8$. The distribution of $\tau_{\rm sc,\,1GHz}$ values obtained from these simulations is shown in Figure 3. We find that a mean value of $\tau_{\rm sc,\,1GHz}=4.13$ ms, or $\tau_{\rm sc,\,600MHz}=31.87$ ms is sufficient to suppress all detections. The maximum $\tau_{\rm sc}$ predicted by Galactic-$n_{e}$ models for the LOS through the center of the detection gap are: $\tau_{\rm sc,\,1GHz}^{\rm NE2001}$ = 0.1 ms and $\tau_{\rm sc,\,1GHz}^{\rm YMW16}$ = 16 ms. This range of model-dependent maximum values of $\tau_{\rm sc}$ is consistent with our inferred maximum derived from the simulations. Moreover, this consistency reinforces the idea that enhanced scattering in Cygnus X is likely to explain the detection gap, though our limit was obtained in a manner independent of either $n_{e}$ model. However, we emphasize that these models are outdated, differ in their treatment of the Cygnus X region [^1] For example, the YMW16 model does not incorporate external information available on the Cyngus X complex., and lack accurate modeling of small-scale plasma fluctuations, so their predictions should be interpreted only as rough references.
5 Discussion
@@
Figure 4: Planck EM map (Planck Collaboration et al. 2016) in Galactic coordinates centered at $l=80\degree$, $b=0\degree$ (top) and $l=180\degree$, $b=0\degree$ (bottom). Overlaid black contours trace logarithmically-spaced EM 10, 100 and 1,000  $\text{ pc cm}^{-6}$. White solid and dashed circles highlight WISE H II regions (L. D. Anderson et al., 2014) with angular radii $\theta_{\text{IR}}$ and $2\theta_{\text{IR}}$, respectively. The radius of the IR emission $\theta_{\text{IR}}$ is provided by the WISE Catalog. H II regions from the HH14 catalog (L. G. Hou & J. L. Han, 2014a) are not shown in this figure as their angular radii are not available but FRBs intersecting them are included (see §5.2 for the treatment). Lime markers are for FRB LOS with intersecting H II regions and skyblue markers are for non-intersecting FRBs. Cross markers indicate FRBs for which no reliable scattering timescale was measured with total-intensity data, while the radii of circular markers are scaled according to the log of the measured scattering timescale.
@@
The statistical significance of the detection gap was aided by its large angular extent, as discussed in §3.1. Smaller spatial gaps however, can occur through chance. We used the non-detections of FRBs in the direction of Cygnus X to place data-driven constraints on the magnitude of maximal scattering from the Milky Way Galaxy, in a manner independent of existing Galactic-$n_{e}$ models. However, even if smaller in angular extents, other ionized structures in the Galaxy such as H II regions and high-EM sightlines may have similar impacts on the observed scattering properties and detectability of FRBs in Catalog 2. This may introduce a selection bias in observed FRB environments, potentially skewing interpretations of their host galaxy associations.
@@
To statistically test whether highly scattered FRBs are under-detected in general due to propagation through regions of elevated emission measure, we compared EMs at FRB positions with those along all sightlines visible to CHIME. Exposure weighting is applied using the CHIME/FRB beam exposure map (CHIME/FRB Collaboration et al. 2025), such that each sky position contributes proportionally to its observing time. Figure 5 shows this comparison using the Planck EM map (Planck Collaboration et al. 2016). The FRB distribution drops to zero for EM $\geq$ 2900 pc cm-6. This difference demonstrates a deficit of FRBs along high EM sightlines that are associated with ionized regions in the Galaxy.
We can ask what the above EM cutoff of 2900 $\text{ pc cm}^{-6}$ corresponds to in terms of $\tau_{ff}$, DM, and $\tau_{\rm sc}$. Similar to the calculations in §4.3, $\tau_{ff}$ is negligible ($\approx 0.003$) in our band. Using Equation 6, and assuming L = 40 pc (a characteristic H II region path length corresponding to typical Strömgren diameter for O8-B0 stars; see §5.2), we estimate a DM contribution near the cutoff of $\approx 340$ $\text{ pc cm}^{-3}$, which is too small to induce appreciable DM smearing. We estimate $\tau_{\rm sc}$ using the empirical $\tau_{\rm sc}-{\rm DM}$ relation from J. M. Cordes et al. (2016), resulting in a rough lower bound of $\tau_{\rm sc,\,1GHz}\gtrsim 0.5\,$ms, or $\tau_{\rm{sc,\,600MHz}}\gtrsim 4\,$ms. We note that use of their relation while assuming DM $\propto\sqrt{\rm EM}$ yields $\tau_{\rm sc}\propto{\rm EM}^{2.25}$ for DM $\gtrsim 100$ $\text{ pc cm}^{-3}$; values of EM above the aforementioned cutoff yield values of $\tau_{\rm sc}$ that rapidly rises to magnitudes that make FRBs undetectable at CHIME. We also note that pulsars behind H II regions lie preferentially above the $\tau_{\rm sc}-{\rm DM}$ relation and that a future reassessment is therefore warranted (S. K. Ocker et al., 2024). While we caution that the logic mapping EM to $\tau_{\rm sc}$ is indirect, the above results support the interpretation that FRBs are preferentially scatter-broadened beyond detectability along high EM sightlines.
@@
5.2 H II Region Intersections and Scattering Properties
Similar to the treatment for pulsar sightlines intersecting H II regions by S. K. Ocker et al. (2024), we utilize 2 complementary catalogs of H II regions – WISE V2.4 (L. D. Anderson et al., 2014) and HH14 (L. G. Hou & J. L. Han, 2014a), as together they span radio, infra-red (IR) and optical wavelengths. The WISE catalog contains over 8,400 H II regions and H II region candidates, identified using the all-sky Wide-field Infrared Survey Explorer (WISE) satellite data. The HH14 catalog contains 4,551 H II regions, of which 611 have spectro-photometric distance measurements. Although HH14 does not provide angular radii of H II regions, we adopt the maximum impact parameter 20 pc from S. K. Ocker et al. (2024) based on typical Strömgren radii for O8-B0 stars and we then scale the angular extents based on the distances of corresponding H II regions provided in the catalog.
We match FRBs to WISE H II regions by computing angular separations in Galactic coordinates and checking whether each FRB falls within twice the angular radius of any H II region ($2\theta_{\rm IR}$). We adopt this maximum impact parameter of $2\theta_{\rm IR}$ following the prescription of S. K. Ocker et al. (2024) to account for potential underestimation of the ionized gas extent by the infrared size, which may miss diffuse outer emission caused by radiation leakage (M. Luisi et al., 2019; J. Dey et al., 2024). Out of the 3552 FRBs in our sample, 36 FRB sightlines pass through WISE H II regions and 85 sightlines pass through HH14 H II regions, with 11 sightlines in common between them, making a total of 110.
Among FRB sightlines that intersect H II regions, as well as those that do not but lie near the Galactic plane ($|b|<10\degree$), only about one-third pass the $\tau_{\rm sc}$ reliability cut off described in §2.4. This sample of sightlines is not sufficiently large to enable a robust statistical comparison of $\tau_{\rm sc}$ between these subsets. For completeness, we report the mean and median $\tau_{\rm sc,\,1GHz}$ values: for intersecting sightlines, the mean is 1.68 ms and the median is 0.63 ms; for non-intersecting near-plane sightlines, the mean is 1.05 ms and the median is 0.53 ms. The intersecting sample exhibits slightly higher values consistent with expectations, but an AD-test between the two populations yields a statistic of 1.58 and a $p$-value of 0.07 indicating no significance. A more robust assessment will require either a larger sample from future CHIME/FRB catalogs or the use of raw baseband data with its higher time resolution to permit measurement of shorter scattering timescales. For interested readers, Table 2 in Appendix B lists the FRB sightlines that intersect H II regions, along with their positions, DMs, and $\tau_{\rm sc}$. These may be useful for further case studies of scattering in dense Galactic environments. Their dynamic spectra are presented in Figure 7 in Appendix B to illustrate observed pulse morphologies.
@@ 5.2 H II Region Intersections and Scattering Properties
We also explore the possibility of a correlation or lack thereof between $\tau_{\rm sc}$ and DM for FRBs. Figure 6 shows a $\tau_{\rm sc}$-DM plot for all FRB candidates in Catalog 2, and exhibits an overall trend that confirms previous findings – FRBs are systematically under-scattered compared to expectations from the $\tau_{\rm sc}$-DM relation from Galactic pulsars (J. Cordes et al., 2016). The discrepancy arises because a large fraction of the FRB DM comes from the intergalactic medium, which contributes minimally to scattering (D. R. Lorimer et al., 2007; D. Thornton et al., 2013). However, this comparison is biased to reflect the statistics of detected events as we likely miss highly scattered FRBs passing through H II regions that therefore fall below the CHIME/FRB detection threshold. For example, Catalog 2 FRBs that intersect Galactic H II regions do not show a strong correlation between DM and $\tau_{\rm sc}$ in Figure 6 for FRBs – even when considering only the Galactic contribution to DM – despite the expectation that significant $\delta n_{e}$ in H II regions causes scattering. However, FRB 20210705 as detected by the FAST telescope (see discussion in §4.1.4) indeed follows the $\tau_{\rm sc}-{\rm DM}$ trend, indicating that its scattering is consistent with being induced by the Cygnus X region. While the statistics currently remain small, this difference raises a cautionary note for population studies: low scattering, high DM bursts may be overrepresented in Catalog 2. CHIME/FRB’s selection function for Catalog 1 has been studied by injecting synthetic FRBs (M. Merryfield et al., 2023). This effort confirmed that the CHIME/FRB backend is more sensitive to bursts with smaller broadened pulse widths. An updated injections study for Catalog 2 is currently underway. In order to make conclusions about the cosmological population of FRBs, a follow-up study incorporating the updated selection functions and host models is essential.
5.3 Visibility Fitting of Scattered FRBs as a Complementary Probe of ISM Scattering
As discussed in 4.1.2, a scattered source will also be angularly broadened, where larger time delays correspond to larger deflection angles. Assuming $d_{\rm eff}\approx d_{\rm scr}$ for an extragalactic source, the mapping between delay $\tau$ and observed angle $\theta$ for each scattered path through a thin screen is (e.g. M. A. Walker et al. 2004)
@@ 5.3 Visibility Fitting of Scattered FRBs as a Complementary Probe of ISM Scattering
Given the mapping between $\tau$ and $\theta$, the interferometric visibility ($V$) of scattered FRBs (or pulsars) will show a characteristic shape as a function of $V(\tau,\nu)$ through a scattering tail (O. Wucknitz, 2012, 2014), which allows for a determination of screen distances and orientations. Conversely, for spatially unresolved sources, an upper limit on the angular broadening will place a lower limit on the screen distance. This will allow mapping of Galactic scattering structures to astrophysical sources, and will cleanly distinguish between Galactic and extragalactic scattering screens.
@@
Upcoming ultra-wideband instruments such as the Canadian Hydrogen Observatory and Radio Transient Detector (CHORD; 300–1500 MHz; K. Vanderlinde et al., 2019) and DSA-2000 (0.7–2GHz; G. Hallinan et al., 2019) will be particularly well-suited to studies of scattering, offering improved sensitivity across a broad frequency range.
6 Conclusion

### v2

@@ IV.5 Empirical Limits on Scattering Timescale
As summarized in §IV.1, the measurements to date of sources observed through the Cygnus X region – through angular broadening of background quasars and Cygnus X-3, scattering of distant pulsars and one FRB detected by FAST – show significant scattering ($\tau_{\rm sc,\,600\,MHz}\sim 10-1000$ ms). This already suggests that scattering can be deleterious for FRB detections at CHIME, and is likely a significant contributor to the lack of CHIME/FRB detections in the gap. However, the magnitude of scattering, which depends on plasma fluctuations on tiny scales at $\ll\rm{AU}$, can drastically differ between sightlines. In this Section, we quantify the effect of scattering on CHIME/FRB detections, and in turn set a lower limit on the average scattering in the detection gap which would lead to zero detections.
@@
We first searched the positional and fitburst data in Catalog 2 for a radial trend in measured $\tau_{\rm sc}$ from the boundary of the Cygnus X region, observing no clear trend. In Catalog 1 injections, CHIME/FRB found a strong selection bias against temporally broad bursts ($W_{b}>10$ ms, calculated at 600 MHz; mts+23). The lack of a trend may be due to the narrow observable window in $\tau_{\rm sc}$: the lower limit set by the reliability cut-off (see §II.4; 0.13 ms at 1 GHz); and the upper limit ($\geq 10$ ms at 1 GHz) that corresponds to scattered pulses likely too broad to detect. Along with a limited number of detected FRBs, a statistically significant radial trend is difficult to establish without finer time resolution to probe a wider range in $\tau_{\rm sc}$. As a result, it is difficult to trace if there is a gradual increase in scattering timescales for sightlines closer to the center of the detection gap.
@@
From the simulations described in §IV.4, we found that on an average 6 bursts are expected to be seen in the detection gap. These simulations account for the elevated sky temperature but not for scattering. Here, we extend the simulation to assess the degree of scattering required to see zero bursts.
@@
Figure 4: Distribution of scattering timescales at 1 GHz that result in zero detectable FRBs in the gap across 106 simulations, incorporating elevated sky temperature as described in §IV.4. Vertical dashed lines indicate scattering predictions from the NE2001 model (black) and YMW16 model (red) for the line of sight along the center of the gap.
@@
In the simulations outlined in §IV.4, we introduce pulse broadening due to scattering using Equation 9. For each simulated realization, we compute the minimum $\tau_{\rm sc}$ using Equation 8, such that the increased $W_{b}$ and $T_{\rm sky}$ reduce the signal-to-noise of all FRBs within the detection gap to $(S/N)_{b}<(S/N)_{\rm thresh}=8$. The distribution of $\tau_{\rm sc,\,1GHz}$ values obtained from these simulations is shown in Figure 4. We find that a mean value of $\tau_{\rm sc,\,1GHz}=5.59$ ms, or $\tau_{\rm sc,\,600MHz}=43.13$ ms is sufficient to suppress all detections. Because all simulations adopt the 99th percentile $T_{\rm sky}$ measured within the void area, these are conservative lower limits on $\tau_{\rm sc}$. Using more typical, lower $T_{\rm sky}$ would require even larger $\tau_{\rm sc}$ to achieve the same suppression in $(S/N)_{b}$. The maximum $\tau_{\rm sc}$ predicted by Galactic-$n_{e}$ models for the LOS through the center of the detection gap are: $\tau_{\rm sc,\,1GHz}^{\rm NE2001}$ = 0.1 ms and $\tau_{\rm sc,\,1GHz}^{\rm YMW16}$ = 16 ms. This range of model-dependent maximum values of $\tau_{\rm sc}$ is consistent with our inferred maximum derived from the simulations. Moreover, this consistency reinforces the idea that enhanced scattering in Cygnus X is likely to explain the detection gap, though our limit was obtained in a manner independent of either $n_{e}$ model. However, we emphasize that these models are outdated, differ in their treatment of the Cygnus X region [^1] For example, the YMW16 model does not incorporate external information available on the Cyngus X complex., and lack accurate modeling of small-scale plasma fluctuations, so their predictions should be interpreted only as rough references.
V Discussion
@@
Figure 5: Planck EM map (planckforegrounds16) in Galactic coordinates centered at $l=80\degree$, $b=0\degree$ (top) and $l=180\degree$, $b=0\degree$ (bottom). Overlaid black contours trace logarithmically-spaced EM 10, 100 and 1,000  $\text{ pc cm}^{-6}$. White solid and dashed circles highlight WISE H II regions (wise) with angular radii $\theta_{\text{IR}}$ and $2\theta_{\text{IR}}$, respectively. The radius of the IR emission $\theta_{\text{IR}}$ is provided by the WISE Catalog. Lime markers are for FRB LOS with intersecting H II regions and skyblue markers are for non-intersecting FRBs. Cross markers indicate FRBs for which no reliable scattering timescale was measured with total-intensity data, while the radii of circular markers are scaled according to the log of the measured scattering timescale.
@@
The statistical significance of the detection gap was aided by its large angular extent, as discussed in §III. Smaller spatial gaps however, can occur through chance. We used the non-detections of FRBs in the direction of Cygnus X to place a data-driven lower limit on the magnitude of maximal scattering contribution from the Milky Way Galaxy, in a manner independent of existing Galactic-$n_{e}$ models. However, even if smaller in angular extents, other ionized structures in the Galaxy such as H II regions and high-EM sightlines may have similar impacts on the observed scattering properties and detectability of FRBs in Catalog 2. This may introduce a selection bias in the observed spatial distribution of FRBs.
@@
To statistically test whether highly scattered FRBs are under-detected in general due to propagation through regions of elevated emission measure, we compared EMs at FRB positions with those along all sightlines visible to CHIME. Exposure weighting is applied using the CHIME/FRB beam exposure map (chimefrbcatalog2), such that each sky position contributes proportionally to its observing time. Figure 6 shows this comparison using the Planck EM map (planckforegrounds16). The FRB distribution drops to zero for EM $\geq$ 2900 pc cm-6. This difference demonstrates a deficit of FRBs along high EM sightlines that are associated with ionized regions in the Galaxy.
We can ask what the above EM cutoff of 2900 $\text{ pc cm}^{-6}$ corresponds to in terms of $\tau_{ff}$, DM, and $\tau_{\rm sc}$. Similar to the calculations in §IV.3, $\tau_{ff}$ is negligible ($\approx 0.003$) in our band. Using Equation 4, and assuming L = 10 pc (a typical H II region path length; see §V.2), we estimate a DM contribution near the cutoff of $\approx 170$ $\text{ pc cm}^{-3}$, which is too small to induce appreciable DM smearing. We estimate $\tau_{\rm sc}$ using the empirical $\tau_{\rm sc}-{\rm DM}$ relation from cordes2016radio, resulting in a rough lower bound of $\tau_{\rm sc,\,1GHz}\gtrsim 0.5\,$ms, or $\tau_{\rm{sc,\,600MHz}}\gtrsim 4\,$ms. We note that use of their relation while assuming DM $\propto\sqrt{\rm EM}$ yields $\tau_{\rm sc}\propto{\rm EM}^{2.25}$ for DM $\gtrsim 100$ $\text{ pc cm}^{-3}$; values of EM above the aforementioned cutoff yield values of $\tau_{\rm sc}$ that rapidly rises to magnitudes that make FRBs undetectable at CHIME. We also note that pulsars behind H II regions lie preferentially above the $\tau_{\rm sc}-{\rm DM}$ relation and that a future reassessment is therefore warranted (OAL+24). While we caution that the logic mapping EM to $\tau_{\rm sc}$ is indirect, the above results support the interpretation that FRBs are preferentially scatter-broadened beyond detectability along high EM sightlines.
@@
V.2 H II Region Intersections and Scattering Properties
Similar to the treatment for pulsar sightlines intersecting H II regions by OAL+24, we utilize the WISE V2.4 catalog of H II regions (wise). The WISE catalog contains over 8,400 H II regions and H II region candidates, identified using the all-sky Wide-field Infrared Survey Explorer (WISE) satellite data. We match FRBs to WISE H II regions by computing angular separations in Galactic coordinates and checking whether each FRB falls within twice the angular radius of any H II region ($2\theta_{\rm IR}$). We adopt this maximum impact parameter of $2\theta_{\rm IR}$ following the prescription of OAL+24 to account for potential underestimation of the ionized gas extent by the infrared size, which may miss diffuse outer emission caused by radiation leakage (luisi+19; dey+24). In Appendix B, we report the angular separations ($\eta$) between H II regions and intersecting FRB sightline candidates, in units of $\theta_{\rm IR}$. Out of the 3542 FRBs in our sample, 36 are candidate sightlines that pass through the WISE H II regions. We similarly examined the optical H II region catalog HH14 (vizierhh14) following the method described in OAL+24. We do not include the results in our main analysis due to the lack of angular radii in HH14, however the results are consistent with those obtained using WISE.
Among candidate FRB sightlines that intersect H II regions, as well as those that do not but lie near the Galactic plane ($|b|<10\degree$), only about one-third pass the $\tau_{\rm sc}$ reliability cut off described in §II.4. This sample of sightlines is not sufficiently large to enable a robust statistical comparison of $\tau_{\rm sc}$ between these subsets. For completeness, we report the mean and median $\tau_{\rm sc,\,1GHz}$ values: for intersecting sightlines, the mean is 1.65 ms and the median is 0.63 ms; for non-intersecting near-plane sightlines, the mean is 1.06 ms and the median is 0.51 ms. The intersecting sample exhibits slightly higher values consistent with expectations, but an AD-test between the two populations yields a statistic of 1.78 and a $p$-value of 0.06 indicating no significance. A more robust assessment will require either a larger sample from future CHIME/FRB catalogs or the use of raw baseband data with its higher time resolution to permit measurement of shorter scattering timescales. For interested readers, Table 2 in Appendix B lists the FRB sightlines that intersect H II regions, along with their positions, DMs, $\tau_{\rm sc}$, and angular separations in units of $\theta_{\rm IR}$. These may be useful for further case studies of scattering in dense Galactic environments. Their dynamic spectra are presented in Figure 8 in Appendix B to illustrate observed pulse morphologies.
@@ V.2 H II Region Intersections and Scattering Properties
We also explore the possibility of a correlation or lack thereof between $\tau_{\rm sc}$ and DM for FRBs. Figure 7 shows a $\tau_{\rm sc}$-DM plot for all FRB candidates in Catalog 2, and exhibits an overall trend that confirms previous findings – FRBs are systematically under-scattered compared to expectations from the $\tau_{\rm sc}$-DM relation from Galactic pulsars (cordes2016radio). The discrepancy arises because a large fraction of the FRB DM comes from the intergalactic medium, which contributes minimally to scattering (lbm+07; tsb+13). However, this comparison is biased to reflect the statistics of detected events as we likely miss highly scattered FRBs passing through H II regions that therefore fall below the CHIME/FRB detection threshold. For example, Catalog 2 FRBs that intersect Galactic H II regions do not show a strong correlation between DM and $\tau_{\rm sc}$ in Figure 7 for FRBs – even when considering only the Galactic contribution to DM – despite the expectation that significant $\delta n_{e}$ in H II regions causes scattering. However, FRB 20210705 as detected by the FAST telescope (see discussion in §IV.1.4) indeed follows the $\tau_{\rm sc}-{\rm DM}$ trend, indicating that its scattering is consistent with being induced by the Cygnus X region. While the statistics currently remain small, this difference raises a cautionary note for population studies: low scattering, high DM bursts may be overrepresented in Catalog 2. CHIME/FRB’s selection function for Catalog 1 has been studied by injecting synthetic FRBs (mts+23). This effort confirmed that the CHIME/FRB backend is more sensitive to bursts with smaller broadened pulse widths. An updated injections study for Catalog 2 is currently underway. In order to make conclusions about the cosmological population of FRBs, a follow-up study incorporating the updated selection functions and host models is essential.
V.3 Visibility Fitting of Scattered FRBs as a Complementary Probe of ISM Scattering
As discussed in IV.1.2, a scattered source will also be angularly broadened, where larger time delays correspond to larger deflection angles. Assuming $d_{\rm eff}\approx d_{\rm scr}$ for an extragalactic source, the mapping between delay $\tau$ and observed angle $\theta$ for each scattered path through a thin screen is (e.g. wmsz04)
@@ V.3 Visibility Fitting of Scattered FRBs as a Complementary Probe of ISM Scattering
Given the mapping between $\tau$ and $\theta$, the interferometric visibility ($V$) of scattered FRBs (or pulsars) will show a characteristic shape as a function of $V(\tau,\nu)$ through a scattering tail (wucknitz12; Wucknitz_GC), which allows for a determination of screen distances and orientations. Conversely, for spatially unresolved sources, an upper limit on the angular broadening will place a lower limit on the screen distance. This will allow mapping of Galactic scattering structures to astrophysical sources, and will cleanly distinguish between Galactic and extragalactic scattering screens.
@@
Upcoming ultra-wideband instruments such as the Canadian Hydrogen Observatory and Radio Transient Detector (CHORD; 300–1500 MHz; vlg+19) and DSA-2000 (0.7–2GHz; hrw+19) will be particularly well-suited to studies of scattering, offering improved sensitivity across a broad frequency range.
VI Conclusion